The activation energy for the reaction NO2(g)+CO(g)⟶NO(g)+CO2(g) is Ea = 175 kJ/mol and the change in enthalpy for the reaction is ΔH = -375 kJ/mol . What is the activation energy for the reverse reaction?
Answer #1

delta H = Eaf - Ear

Eaf is activation energy for forward direction

Ear is activation energy for reverse direction

so,

delta H = Eaf - Ear

-375 = 175 - Ear

Ear = 175 + 375

= 550 KJ/mol

Answer: 550 KJ/mol